Q1. Martha is preparing for exams in economics and sociology. She has time

to read 40 pages of economics and 30 pages of sociology. In the same

amount of time, she could also read 30 pages of economics and 60 pages of

sociology.

(a) Assuming that the number of pages per hour that she can read of either

subject does not depend on how she allocates her time, how many pages

of sociology could she read if she decided to spend all of her time on

sociology and none on economics?

(b) How many pages of economics could she read if she decided to spend

all of her time reading economics?


1
Expert's answer
2021-11-15T10:16:09-0500

Solution:

Q1.) a.). Budget bundles (40, 30) and (30, 60)

y = ax + b

Slope of the budget line = Δy"\div" Δx = -30"\div" 10 = -3

Utilizing the first bundle: (40, 30)

30 = -3 x 40 + b

30 + 120 = b

b = 150

If all her time is spent on sociology and none on economics:

When x = 0, y = 150

She will read 150 pages of sociology

 

b.). y = ax + b

If all her is spent on reading economics:

When y = 0, then x = 50

She will read 50 pages of economics